OpenStudy (heyitszia1):

Please help, fan and medal! Create a system of equations that includes one linear equation and one quadratic equation. Part 1. Show all work to solving your system of equations algebraically. Part 2. Graph your system of equations, and show the solution graphically to verify your solution.

OpenStudy (heyitszia1):

@jigglypuff314 @imqwerty @NekoMatt

OpenStudy (heyitszia1):

@myininaya @Needhelpp101 @sleepyjess

OpenStudy (agent0smith):

Use y, not f(x) or g(x).

OpenStudy (agent0smith):

Now to solve, notice they're both y = which means you can set them equal to each other.

OpenStudy (agent0smith):

Yep, now get it equal to zero, get all the terms on one side. Then you can either factor or use quadratic formula.

OpenStudy (heyitszia1):

@agent0smith \[0 = 2x^2\]

OpenStudy (will.h):

Correct so far

OpenStudy (will.h):

That was for part A so part B you can use a graph system like GeoGebra to plug the equation and you'll get the graph.

OpenStudy (agent0smith):

Now solve for x.

OpenStudy (agent0smith):

Then once you have x go back and find y

OpenStudy (agent0smith):

@HeyItsZia1

OpenStudy (agent0smith):

No. Get x^2 alone FIRST.

OpenStudy (heyitszia1):

Okay. \[0=2x^2\] \[0\div2=x^2\] \[0=x^2\] ?

OpenStudy (heyitszia1):

@agent0smith Is this right?

OpenStudy (agent0smith):

Yes. Now you can square root.

OpenStudy (heyitszia1):

\[\sqrt{0}=x\] So x would be 0?

OpenStudy (agent0smith):

Yes.

OpenStudy (agent0smith):

That doesn't help you solve for y...

OpenStudy (heyitszia1):

y = 24

OpenStudy (heyitszia1):

For part B, is this okay? https://www.desmos.com/calculator/bhuirwczfa

OpenStudy (heyitszia1):

@agent0smith

OpenStudy (agent0smith):

Looks fine